B. Find the mean (𝝁) variance (𝝈 2)and standard deviation (𝝈) of the population given the sample size, and the means and variances of the sampling distribution of the means. Sample size n were randomly selected from the population



4. 𝑛 = 5

πœ‡π‘₯ = 3

𝜎²π‘₯ = 4.5

Mean:


Variance:


Standard Deviation:


5. 𝑛 = 7

πœ‡π‘₯ = 10

𝜎²π‘₯ = 6

Mean:


Variance:


Standard Deviation:


1
Expert's answer
2022-04-04T16:42:24-0400

Mean of the population is equal to the mean of the sampling distribution, variance is equal to sample distribution variance multiplied by n. So,

1) "\mu=\mu_x=3"

"\sigma^2=\sigma^2_x*5=4.5*5=22.5"

"\sigma=\sqrt{\sigma}=\sqrt{22.5}\approx4.74"


2) "\mu=\mu_x=10"

"\sigma^2=\sigma^2_x*7=6*7=42"

"\sigma=\sqrt{\sigma}=\sqrt{42}\approx6.48"